Solve.
√(x-1)-√(15x-1)=0
The solution is x= ____.

Answer by lwsshak3(11628)

Solve.
√(x-1)-√(15x-1)=0
...
√(x-1)-√(15x-1)=0
√(x-1)=√(15x-1)
square both sides
x-1=15x-1
x≠15x
no solution